OpenStudy (anonymous):

a particle moves along x axis in such a way that its x-coordinates varies with time t according to the relation x-2- 5t +6t^2 the imitial velocity (at t=0) of the particle is?

OpenStudy (anonymous):

do you mean x=2- 5t +6t^2 instead of x-2- 5t +6t^2 that minus, is it right? or...?

OpenStudy (anonymous):

the ans=-5 units

OpenStudy (anonymous):

can u solve for it

OpenStudy (anonymous):

well i think all have to be in terms of t BUt i'd take the derivitive of x(t) then evaluate it at 0 x'(0)

OpenStudy (anonymous):

hmm if x is correct then i'd use implicit differentiation x=x(t)-2- 5t +6t^2

OpenStudy (anonymous):

i think should be the relation x= 2-5t + 6t^2 v=dx/dt = -5 + 12t so when t=0 v= -5

OpenStudy (anonymous):

thanzzzz

OpenStudy (anonymous):

yeah it is easy when x=2-5t...